# Revisiting Graph-based Fraud Detection in Sight of Heterophily and Spectrum

### v1

@@ Introduction
BaseLines for Comparison. We conduct comparative experiments between SEC-GFD and advanced baselines that belong to three groups. The first group consists of methods based on homophily GNN models, including GCN(Kipf and Welling 2016), GraphSAGE(Hamilton, Ying, and Leskovec 2017), GAT(Veličković et al. 2017) and GIN(Xu et al. 2018). The second group comprises state-of-the-art methods for graph-based fraud detection, namely GraphConsis(Liu et al. 2020), CARE-GNN(Dou et al. 2020), PC-GNN(Liu et al. 2021b) and GAGA(Wang et al. 2023). The third group contains GNN-based fraud detection models specifically designed for heterophily, including ACM(Luan et al. 2022), H2-FDetector(Shi et al. 2022), BWGNN(Tang et al. 2022) and GDN(Gao et al. 2023b). More details on description and implementation of the benchmarks are provided in Appendix B.
Metrics. We choose two widely used metrics to measure the performance of all the methods, namely F1-macro and AUC. F1-macro is the macro average of F1 score, which neglects the imbalance ratio between normal and anomaly labels, and AUC is the area under the ROC Curve.
Settings. In our proposed method SEC-GFD, for the Amazon, YelpChi, and T-Finance datasets, the hidden layer dimension is set to 64, and the high-frequency signal neighbor order C is set to 2. For the T-Social dataset, the hidden layer dimension is set to 16, the order C of high-frequency signal neighbors is set to 5. The experimental results show that, for such a large-scale graph, fusing high-frequency information of higher-order neighbors can learn better features. In the experiment, to ensure fairness, the size of training/validation/testing set of the datasets is set to 0.4/0.2/0.4 for all the compared methods. Our method is implemented with the DGL library in Pytorch, and for other methods, public source codes are utilized for implementation. Furthermore, we run for 100 epochs on the four datasets for all methods.
@@
We conduct a comprehensive comparison of our proposal with vanilla homophily GNN models, state-of-the-art graph-based fraud detection models and novel GNNs targeted at heterophily. The corresponding results are shown in Tab. 2.
First of all, It is observed that SEC-GNN outperforms all vanilla homophily GNN models, i.e., GCN, GraphSAGE, GAT and GIN. For instance, when compared with GIN, SEC-GFD achieves improvements of 13.37% and 26.41% in terms of F1-macro and AUC on Amazon. To delve reasons, these models can be regarded as low-pass filters in essence and only focus on low frequency information. While our SEC-GNN integrates rich information of distinct frequency bands to adapt to heterophily.
When compared with state-of-the-art graph-based fraud detection methods, these methods still have a large gap with our proposal. Among them, GraphConsis has the worst comprehensive performance. It combines the characteristics of nodes with the surrounding environments to learn a relationship attention weight, which is essentially a low-pass filter. CARE-GNN and PC-GNN achieve some improvements in the effect. This is because they are aware of heterophily edges and ”camouflage” problems, and design corresponding neighbor selection module. GAGA introduces a group aggregation module to generate distinguishable multi-hop neighborhood information. Although it demonstrates substantial overall improvement, it still exhibits a significant disparity when compared to our proposed method. Specically, when compared with GAGA, SEC-GFD achieves gains of 2.04% and 2.62% in terms of F1-macro and AUC on YelpChi. The reason behind this can be attributed to that our method additionally learns complicated high-frequency information in addition to learning multi-hop neighbor information.
Finally, when compared with novel heterophily GNNs, experimental results in Tab. 2 show that the proposed method consistently outperforms them on all metrics across the four datasets. To be specific, SEC-GFD attains improvements of 2.47% and 1.71% in terms of F1-macro and AUC on T-Finance. As to T-Social, SEC-GFD achieves gains of 3.78% and 0.84% in terms of F1-macro and AUC. Among these models, ACM studies heterophily in terms of feature similarity, H2-FDetector identifies homophily and heterophily connections and applies different aggregation strategies, BWGNN designs a band-pass filter to transmit information in multiple frequency bands, and GDN divides the node features into anomaly pattern and neighborhood pattern. The superior performance of SEC-GFD indicates the effectiveness of the proposed hybrid filtering module and local environmental constraint module.
@@
To explore the effect of heterophily edge pruning on GNNs with different filter properties, we conduct extensive experiments. In this paper, three message passing mechanisms with low-pass, high-pass, and band-pass filtering characteristics are designed for experiments. The low-pass and high-pass filters use the adjacency matrix and normalized Laplacian matrix respectively for message passing, and the band-pass filter uses the beta wavelet filter function from BWGNN, with two layers of neighbor information are considered. The experimental results on Amazon and Yelpchi are in Fig. 3.
Figure 3: Filters’ performance with different heterophily edges’ deletion ratio on Amazon and YelpChi. Fig(a) and Fig(c) denotes deletion of heterophily edges on the entire graph, while Fig(b) and Fig(d) denotes deletion of heterophily edges appearing in training graph.
Considering the difficulty in discriminating heterophily edges, in many cases only some heterophily edges connected with labeled nodes on the training set can be identified. Thus two sets of experiments are conducted in this paper. As observed in Fig. 3, when clipping the heterophily edges of the whole graph, the performance of low-pass and band-pass filtering improves with the ratio of heterophily edge clipping. While the performance of high-pass filtering shows a monotonous decrease. And the performance of the band-pass filter keeps the best. Moreover, when clipping heterophily edges appearing in the training set, results are completely different. Specifically, low-pass filter maintains a poor performance, while band-pass and high-pass filter exhibit different degrees of declining trends. This suggests that indiscriminate removal of heterophily edges may not necessarily be beneficial.
@@
To explore how each component contributes to the performance of SEC-GFD, we systematically remove certain parts of the framework for ablation study. The results are shown in Tab. 3.
Firstly, in order to explore the performance of the two parts of the hybrid filter, we omit band-pass bands and low-order pure high-pass bands respectively. When band-pass bands are excluded, the performance drops by 2.06% and 2.67% in terms of AUC and F1-macro. And the performance decreases 1.13% and 1.27% when low-order high-pass bands are removed in terms of AUC and F1-macro. It is obviously that band-pass bands are more important, this is because band-pass bands consider high frequency and lower frequency information together. Afterwards, we delete the local environmental constraint, the performance drops by 0.86% and 1.31% in terms of AUC and F1-macro. However, it still outperforms other state-of-the-art methods, this also reflects the effectiveness of our hybrid filter. Thus the hybrid filter can be equipped as a general backbone to various other GNN-based fraud detectors to improve performance.
Table 3: Ablation Performance
@@
F1-macro
AUC
F1-macro
AUC
SEC-GFD
92.27 98.19 89.74 96.34
w/o high-pass
91.48
97.28 88.65
95.05
w/o band-pass
90.29
96.31
86.79
93.77
w/o env
91.88
97.14
88.14
95.21
@@
In the previous article, based on the formula of beta wavelet, we judged that C represents the order of neighbor information considered in the network training process. Therefore, the order C is a very important parameter. Fig. 4 shows the changes of C from 1 to 6 on four real data sets, and shows the changes of F1 and AUC values.
Figure 4: Performance with different order C
@@ Conclusion
This paper reviews the research on highly heterophily graphs, especially in fraud detection scenarios, and further explores the association between heterophily edges and spectral domain energy distribution. The experimental results prove that it is not necessary to discriminate and cut heterophily edges, but to make full use of each frequency band for information transmission to achieve excellent results. To this end, we design a spectral-domain filter-based semi-supervised graph neural network fraud detector SEC-GFD. The main body is a hybrid filtering module, which contains multiple frequency bands with multi-order neighbors. Then it is supplemented with a local environmental constraint module to improve the utilization of node labels and the adaptive relationship between nodes and their surrounding environments.
@@ Acknowledgments
This research is sponsored in part by the NSFC Program (No. U20A6003, 62076146, 62021002, U19A2062, U1911401, 6212780016) and Industrial Technology Infrastructure Public Service Platform Project ”Public Service Platform for Urban Rail Transit Equipment Signal System Testing and Safety Evaluation” (No. 2022-233-225), Ministry of Industry and Information Technology of China.

### v3

@@ Introduction
Table 2: Fraud detection performance on four real-world datasets in terms of F1-macro and AUC value. (The bold and italicized values in the table represent the best and second-best performance, respectively.)
BaseLines for Comparison. We conduct comparative experiments between SEC-GFD and advanced baselines that belong to three groups. The first group consists of methods based on homophily GNN models, including GCN (Kipf and Welling 2016), GraphSAGE (Hamilton, Ying, and Leskovec 2017), GAT (Velivcković et al. 2017) and GIN (Xu et al. 2018). The second group comprises state-of-the-art methods for graph-based fraud detection, namely GraphConsis (Liu et al. 2020), CARE-GNN (Dou et al. 2020), PC-GNN (Liu et al. 2021b) and GAGA (Wang et al. 2023). The third group contains GNN-based fraud detection models specifically designed for heterophily, including ACM (Luan et al. 2022), H2-FDetector (Shi et al. 2022), BWGNN (Tang et al. 2022) and GDN (Gao et al. 2023b). More details on description and implementation of the benchmarks are provided in Appendix B.
Metrics. As graph anomaly detection poses a class-imbalanced classification problem, this paper utilizes two widely adopted metrics: F1-macro and AUC. F1-macro considers the weighted average of F1 scores across multiple classes, and AUC is the area under the ROC Curve.
Settings. In our proposed method SEC-GFD, for the Amazon, YelpChi, and T-Finance datasets, the hidden layer dimension is set to 64, and the high-frequency signal neighbor order C is set to 2. For the T-Social dataset, the hidden layer dimension is set to 16, the order C of high-frequency signal neighbors is set to 5. The experimental results show that, for such a large-scale graph, fusing high-frequency information of higher-order neighbors can learn better features. In the experiments, to ensure fairness, the size of training/validation/testing set of the datasets is set to 0.4/0.2/0.4 for all the compared methods. Our method is implemented with the DGL library in Pytorch, and for other methods, public source codes are utilized for implementation. Furthermore, we run for 100 epochs on the four datasets for all methods.
@@
We conduct a comprehensive comparison of our approach with vanilla homophily GNN models, state-of-the-art graph-based fraud detection models and novel GNNs targeted at heterophily. The corresponding results are shown in Table 2.
First of all, It is observed that SEC-GNN outperforms all vanilla homophily GNN models, i.e., GCN, GraphSAGE, GAT and GIN. For instance, when compared with GIN, SEC-GFD achieves improvements of 13.37% and 26.41% in terms of F1-macro and AUC on Amazon. To delve reasons, these models can be regarded as low-pass filters in essence and only focus on low frequency information. While our SEC-GNN integrates rich information of distinct frequency bands adapting to heterophily.
When compared with state-of-the-art graph-based fraud detection methods, these methods still have a large gap with our approach. Among them, GraphConsis demonstrates the worst comprehensive performance. It combines the characteristics of nodes with the surrounding environments to learn relationship attention weights, and it is essentially a low-pass filter. CARE-GNN and PC-GNN achieve some improvements in the effectiveness. This is because they are aware of heterophily edges and camouflage problems, and design corresponding neighbor selection module. GAGA introduces a group aggregation module to generate distinguishable multi-hop neighborhood information. Although it demonstrates substantial overall improvements, it still exhibits a significant disparity when compared to our proposed method. Specically, when compared with GAGA, SEC-GFD achieves gains of 2.04% and 2.62% in terms of F1-macro and AUC on YelpChi. The reason behind this can be attributed to that our method additionally learns complicated high-frequency information in addition to learning multi-hop neighbor information.
Finally, when compared with novel heterophily GNNs, experimental results in Table 2 show that our proposed method consistently outperforms them on all metrics across the four datasets. To be specific, SEC-GFD attains improvements of 2.47% and 1.71% in terms of F1-macro and AUC on T-Finance. As to T-Social, SEC-GFD achieves gains of 3.78% and 0.84% in terms of F1-macro and AUC. Among these models, ACM studies heterophily in terms of feature similarity, H2-FDetector identifies homophily and heterophily connections and applies different aggregation strategies, BWGNN designs a band-pass filter to transmit information in multiple frequency bands, and GDN divides the node features into anomaly pattern and neighborhood pattern. The superior performance of SEC-GFD indicates the effectiveness of the proposed hybrid filtering module and local environmental constraint module.
@@
To explore the effect of heterophily edge pruning on GNNs with different filter properties, we conduct extensive experiments. In this paper, three message passing mechanisms with low-pass, high-pass, and band-pass filtering characteristics are designed for experiments. The low-pass and high-pass filters use the adjacency matrix and normalized Laplacian matrix respectively for message passing, and the band-pass filter uses the beta wavelet filter function from BWGNN, with two layers of neighbor information are considered. The experimental results on Amazon and Yelpchi are illustrated in Figure 3.
Figure 3: Filters’ performance with different heterophily edges’ deletion ratio on Amazon and YelpChi. Fig (a) and Fig (c) denote deletion of heterophily edges on the entire graph, while Fig (b) and Fig (d) denote deletion of heterophily edges appearing in training graph.
Considering the difficulty in discriminating heterophily edges, in many cases only some heterophily edges connected with labeled nodes on the training set can be identified. Thus two sets of experiments are conducted in this paper. As observed in Figure 3, when clipping the heterophily edges of the whole graph, the performance of low-pass and band-pass filtering improves with the ratio of heterophily edge clipping. While the performance of high-pass filtering shows a monotonous decrease. And the performance of the band-pass filter keeps the best. Moreover, when clipping heterophily edges appearing in the training set, results are completely different. Specifically, low-pass filter maintains a poor performance, while band-pass and high-pass filter exhibit different degrees of declining trends. This suggests that indiscriminate removal of heterophily edges is not necessarily beneficial.
@@
To explore how each component contributes to the performance of SEC-GFD, we systematically remove certain parts of the framework for ablation study. The results are illustrated in Table 3.
Firstly, in order to explore the performance of the two parts of the hybrid filter, we omit band-pass bands and low-order pure high-pass bands respectively. When band-pass bands are excluded, the performance drops by 2.06% and 2.67% in terms of AUC and F1-macro. And the performance decreases 1.13% and 1.27% in terms of AUC and F1-macro when low-order high-pass bands are removed. It is obviously that band-pass bands are more important, this is because band-pass bands consider high frequency and lower frequency information together. Afterwards, we delete the local environmental constraint, then performance drops by 0.86% and 1.31% in terms of AUC and F1-macro. However, it still outperforms other state-of-the-art methods, this also reflects the effectiveness of our hybrid filter. Thus the hybrid filter can be equipped as a general backbone to various other GNN-based fraud detectors to improve performance.
@@
F1-macro AUC F1-macro AUC
SEC-GFD 92.27 98.19 89.74 96.34
w/o high-pass 91.48 97.28 88.65 95.05
w/o band-pass 90.29 96.31 86.79 93.77
w/o env 91.88 97.14 88.14 95.21
Table 3: Ablation Performance.(The bold and italicized values denote the best and runner-up performance.)
@@
In the previous discussion, based on the formula of beta wavelet, we judge that C represents the order of neighbor information considered in the network training process. Therefore, the order C is an extremely important parameter. Figure 4 shows the changes of F1 and AUC values with the changes of C from 1 to 6 on four real-world datasets.
Figure 4: Performance with different order C.
@@ Conclusion
This paper reviews the research on highly heterophily graphs, especially in fraud detection scenarios, and further explores the association between heterophily edges and spectral domain energy distribution. The experimental results prove that it is not necessary to discriminate and cut heterophily edges, but to make full use of each frequency band for information transmission to achieve excellent results. To this end, we have designed a spectrum-enhanced and environment-constrainted graph-based fraud detector SEC-GFD. The main body is a hybrid filtering module, which contains multiple frequency bands with multi-order neighbors. Then it is supplemented with a local environmental constraint module to improve the utilization of node labels and the adaptive relationships between nodes and their surrounding environments.
@@ Acknowledgments
This work is supported in part by the Fundamental Research Funds for the Central Universities under Grant 2022RC026; in part by the CCF-NSFOCUS Open Fund; in part by the NSFC Program under Grant 62202042, Grant U20A6003, Grant 62076146, Grant 62021002, Grant U19A2062, Grant 62127803, Grant U1911401 and Grant 6212780016; in part by Industrial Technology Infrastructure Public Service Platform Project ‘Public Service Platform for Urban Rail Transit Equipment Signal System Testing and Safety Evaluation’ (No. 2022-233-225), Ministry of Industry and Information Technology of China.
Appendix A A. Bond between Rayleigh Quotient and Spectral Energy
For the Rayleigh quotient, we initially examine its numerator, which corresponds to the dot product of the signal with itself following the Laplacian matrix transformation. Subsequently, we proceed to expand the formula and obtain:
$\begin{split}X^{T}LX&=X^{T}(D-A)X=X^{T}DX-X^{T}AX\\ &=\sum_{i=1}^{N}x_{i}^{2}d_{i}-\sum_{i=1}^{N}\sum_{j=1}^{N}A_{ij}x_{i}x_{j}\\ &=\frac{1}{2}(\sum_{i=1}^{N}x_{i}^{2}d_{i}-2\sum_{i=1}^{N}\sum_{j=1}^{N}A_{ij}x_{i}x_{j}+\sum_{j=1}^{N}x_{j}^{2}d_{j})\\ &=\frac{1}{2}\sum_{i=1}^{N}\sum_{j=1}^{N}A_{ij}(x_{i}-x_{j})^{2}\end{split}$
Upon observation, we note that the numerator is similar to a weighted sum of Euclidean distances, which can also serve as a metric for signal dissimilarity. Subsequently, we substitute this into the formula of the Rayleigh quotient:
$\begin{split}\noindent R(L,X)=\frac{X^{T}LX}{X^{T}X}&=\frac{\sum_{i=1}^{N}\sum_{j=1}^{N}A_{ij}(x_{i}-x_{j})^{2}}{2\sum_{i=1}^{N}x_{i}^{2}}\\ &=\frac{\sum_{(i,j)\in\mathcal{E}}(x_{i}-x_{j})^{2}}{2\sum_{i\in\nu}x_{i}^{2}}\end{split}$
Thus Rayleigh quotient is actually the standardized variance fraction of signal X. To establish a correlation with spectral energy, the analysis is conducted in the spectral domain. According to spectral graph theory, $X^{T}LX=\sum_{i=1}^{N}\lambda_{i}\tilde{x}_{i}^{2}$ and $X^{T}X=\sum_{i=1}^{N}x_{i}^{2}$ have been empirically validated, thus the Rayleigh quotient can also be expanded as follows:
$\begin{split}R(L,X)&=\frac{X^{T}LX}{X^{T}X}=\frac{\sum_{i=1}^{N}\lambda_{i}\tilde{x}_{i}^{2}}{\sum_{i=1}^{N}x_{i}^{2}}\\ &=\sum_{i=1}^{N}\lambda_{i}(\frac{\tilde{x}_{i}^{2}}{\sum_{k=1}^{N}\tilde{x}_{k}^{2}})\\ &=\sum_{i=1}^{N}\lambda_{i}(\eta_{i}-\eta_{i-1})\\ &=\lambda_{N}-\left[(\lambda_{1}-\lambda_{0})\eta_{0}+\cdots+(\lambda_{N}-\lambda_{N-1})\eta_{N-1}\right]\\ &=\lambda_{N}-\left[\int_{0}^{\lambda_{1}}f(t)d_{t}+\cdots+\int_{\lambda_{N-1}}^{\lambda_{N}}f(t)d_{t}\right]\\ &=\lambda_{N}*f(t)_{max}-S_{spec}\end{split}$
where $S_{spec}$ is the area under spectrum energy accumulation curve and $\lambda_{N}$ represents the total length of the continuous spectrum. Consequently, the Rayleigh quotient exhibits a higher value when the proportion of high-frequency components is greater.
Appendix B B. Baselines and Implementation Details
Baselines can be classified into three distinct categories. The first category contains GNNs based on homophily assumption:
  • GCN (Kipf and Welling 2016): GCN is a traditional graph neural network model that aggregates neighbor information.
  • GraphSAGE (Hamilton, Ying, and Leskovec 2017): GraphSAGE is a graph neural network model that samples and aggregates information from node neighborhoods to learn node representations.
  • GAT (Velickovic et al. 2017): GAT is a graph neural network model that applies attention mechanisms to process graph data.
  • GIN (Xu et al. 2018): GIN is a graph neural network model that aggregates node information using permutation-invariant operations.
The second category are state-of-the-art graph-based fraud detection methods:
  • GraphConsis (Liu et al. 2020): GraphConsis combines the characteristics of nodes with the surrounding environment to learn a relationship attention weight.
  • CARE-GNN (Dou et al. 2020): CARE-GNN achieves the effect of modifying the adjacency matrix and pruning the edges by designing the neighbor selection module.
  • PC-GNN (Liu et al. 2021): PC-GNN introduces a node sampler and a label-aware neighbor selector to further reweight imbalanced classes.
  • GAGA (Wang et al. 2023): GAGA introduces a group aggregation module to generate distinguishable multi-hop neighborhood information.
The third category consists of GNN methods specifically designed for heterophily:
  • ACM (Luan et al. 2022): ACM uses aggregated features to study heterophily in terms of feature similarity.
  • H2-FDetector (Shi et al. 2022): H2-FDetector identifies homophily and heterophily connections and applies different aggregation strategies.
  • BWGNN (Tang et al. 2022): BWGNN designs a band-pass filter to transmit information in multiple frequency bands.
  • GDN (Gao et al. 2023): GDN divides the node features into two parts that describe the abnormal pattern and adapt to the neighborhood information.